Suppose that we flip 7 fair coins - what is the probability that we have at least two tails showing?
Answer #1

X ~ Bin ( n,p)

Where n = 7 , p = 0.50 (Probability of head and tail is equally likely, that is 1/2)

Binomial probability distribution is

P(X) = nCx * px * ( 1 - p)n-x

P(X >= 2) = 1 - P(X <= 1)

= 1 - [ P(X = 0) + P(X = 1) ]

= 1 - [ 7C0 * 0.500 * 0.507 + 7C1 * 0.501 * 0.506 ]

= 1 - 0.0625

= 0.9375